OpenStudy (theislander):

Given tan theta=4/3 and sin(-4/5), find cos theta

OpenStudy (theislander):

answer choices

hartnn (hartnn):

use the identity \(\tan \theta = \dfrac{\sin \theta }{\cos \theta} \\ \implies \cos \theta = \dfrac{\sin \theta}{\tan \theta}\)

OpenStudy (theislander):

I got -3/5. Is that right?

hartnn (hartnn):

absolutely! :)

OpenStudy (theislander):

thank you!